OpenStudy (anonymous):

1/2(2)+-(1) can be simplified to 1+(-1) . Name the property that was used to reach this step. A. Multiplicative Inverse Property B. Associative Property of Multiplication C. Multiplicative Identity Property D. Distributive Property

OpenStudy (anonymous):

E ? really...lol

OpenStudy (anonymous):

@massivewand unnecessary...

OpenStudy (anonymous):

goodbye!

OpenStudy (accessdenied):

It appears that the simplification made between the first and second expression was that 1/2(2) changed to 1 That eliminates the Associative property (a*(b*c) = (a*b)*c) and Distributive Property. (a(b+c) = ab + ac) as answer choices. Multiplicative Identity property is that anything multiplied by 1 = itself That leaves one answer remaining. :) Multiplicative Inverse property says that a * 1/a = 1 for any nonzero value replacing a; 5/5 = 1, -2/(-2) = 1, etc.

OpenStudy (accessdenied):

* Which is what we can see in the simplification, that 1/2 * 2 = 1; 2 and its inverse 1/2 multiply together to get 1.